Instrument Log

 

CO2 Flux / Sampling System†††† CF†† Lamont, Oklahoma††††† 60M Tower

 

Date

Technician

Comments

5/31/2016

MS

Noticed that during the trimming of the crop below and around the instrument that the CO2 Flux had rotated from facing down which that is the original position it should be in, but it was facing sideways (East).Trimmed crop below and around instrument upon mentors request.Rotated and tightened down instrument back to its original positing which is facing down.RTS at 1945 GMT.

5/17/2016

MG

Per BCR-02040 Scheduled Patch Release. Re booted the Co2 flux computer and had to power cycle the 4 M system RTS 1450 GMT.

5/6/2016

MG

Husker Flux Programs stoped running the husker flux programs were intermittent throughout the night. Mentor did some programming work and I power cycled all 3 systems RTS 1630 GMT.

4/19/16

MG

Instrument needs to be shut down and restarted due to a scheduled Patch Release at 1300 GMT. At 1609 GMT I shutdown and reset computer to install a Patch as per BCR-02040.RTS at 1614 GMT.

4/7/16

Mentor

Instrument mentor on site to install new soil probes on a smaller, single pole to replace the larger swing set type frame.The old probes and frame were left in place for now, but will be removed later this month.Installed new datalogger and soil probes as listed below.New setup operational at 1800 GMT.

 

- Datalogger (Mfr: Campbell Scientific, Model: CR1000, S/N: 44331, Calibration date: January 27, 2016), Calibration interval: 3 years
- Soil heat flux plate (Mfr: Huskeflux, Model: HFP01SC-15, S/N: 4183)
- Soil heat flux plate (Mfr: Huskeflux, Model: HFP01SC-15, S/N: 4186)
- Soil heat flux plate (Mfr: Huskeflux, Model: HFP01SC-15, S/N: 4187)
- Soil heat flux plate (Mfr: Huskeflux, Model: HFP01SC-15, S/N: 4188)
All sensors were calibrated on October 13, 2014

6 Soil Moisture & Temperature sensors (Mfr: Campbell Scientific, Model: CS655)

3/15/16

MG

Due to a scheduled Patch Release the CO2 Flux needed to be restarted. CO2 Flux 60 M starts on its own no problems. CO2 Flux 25 M and 4 M will not restart with out a full hard power cycle. CO2 Flux 25M First thing I do is close the HuskerFlux 25M Program. Go out to the tower center box on the north side of the tower. Open it and turn the "lower B" switch off. Then go start the HuskerFlux 25M program while the switch is turned off. Obviously it will not start. Once it tells you that it cannot run then close the HuskerFlux 25M program and go turn the switch back on at the tower.Then go start the HuskerFlux 25M program. For the CO2 Flux 4M system close the HuskerFlux 4M program. Unplug the orange plug on the inside the south end of the building. The Power strip says UPS Power on it and I wrote 4M on the plug. Leave unplugged for several minutes then plug back in and restart the program. RTS all CO2 Flux Systems 1857.

2/16/16

MG

Due to IT department computer needed rebooted for patch releases Rebooted the CO2MET Computer at 15:06 RTS 15:24 GMT

2/16/16

MG

Due to IT department computer needed rebooted for patch releases Rebooted Computer at 15:06 the 25 meter system was the only one that came back on its own. I had to power cycle the 60 m switch on the tower base then restart the program. The 4 M system is very temperamental I unplugged the orange cord at the north end of the building then start the huskerflux 4m it will not launch then plug the cord back in and try husker flux 4m again.RTS 15:27RTS 15:27†† GMT

1/19/16

MG

IT department installed software patches on the instrument computer and the system needed to be rebooted. At 20:28 on 1-19-2016 I restarted the CO2 MET computer RTS 2101 GMT

1/19/16

MG

Heat has been hard to regulate due to lack of heater in PGS Building. I bought 2 Osculating heaters and installed a fan to circulate air evenly. This is a temporary fix until new unit gets installed.

1/19/16

MG

IT department installed software patches on the instrument computer and the system needed to be rebooted. At 20:27 on 1-19-2016 I restarted the CO2 Flux computers and restarted the Husker Flux programs RTS 2101 GMT.

12/31/15

MG

The Central Facility lost power due to an ice storm on Sunday 12-27-2015 around 00:00 GMT I reset the building UPS and started program RTS 12-31-2015 at 15:45 GMT.

12/15/15

MG

CO2 Met computer needed to be restarted following scheduled software and security patch updates. Rebooted system at 1916 GMT. RTS 1929 GMT.

12/15/15

MG

Computer needed to be restarted following scheduled software and security patch updates. Rebooted system at 1856 GMT. RTS 1915 GMT.

11/17/15

MG

SGP experienced a power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/15. Instrument was down. No data being displayed for 25m system. Power cycled switch on tower and restarted husker flux program. all is well on 25 M flux system. RTS 1959

11/16/2015

MS

No data being displayed for 60m system.Tried to power cycle instrument but it did not work, at 1815 GMT lowered the west carriages slightly and re-raised into position to re-seat connection at 1819 GMT.Restarted data collection at 1830 and graphs are now displayed.RTS 1830 GMT.

11/13/15

MG

60 M CO2 Flux was not running. Reset the switch on the tower for the 60 M restarted the HuskerFlux program and all is well RTS 1918 GMT

11/12/15

MG

60 M CO2 Flux was not running. Reset the switch on the tower for the 60 M restarted the HuskerFlux program and all is well RTS 1800 GMT

10/28/14

CAM

Lowered the West booms at 1454 GMT for T/H calibration changes.Raised booms and RTS at 2012 GMT.No problem with this instrument.

10/15/2015

MS

Program for the 25m, and the 60m tower was not running.Restarted program at 1616 GMT, RTS at 1631 GMT.

10/2/15

MG

I tried to log in and failed system said user name and password were incorrect. Spoke with David Swank from our IT department he then was able to log in and I restarted the programs all is good.

9/18/15

MDG

Rebooted the computer and the 60M program still would not start collecting. I then power cycled the west 60M level via switches at tower base. Husker-flux program took off.

9/10/15

PD

4M CO2 HuskerFlux program has stopped.Exited/restarted program but it failed to start correctly.Powercycled 4M system and was able to start program.RTS at 1507 GMT.

8/18/15

PD

60M Sonic:4M and 25M systems showed errors and attempts failed to restart program.Stopped 60M program to reboot computer and attempt to restart all 3 programs.Stopped program at 1415 GMT and rebooted computer.4 and 25M program started correctly, 60M program did not.Powercycled 60M instrument but program still failed to restart.Powercycled again and program restarted correctly.RTS at 1430 GMT.

8/18/15

PD

25M Sonic:Error shown:Program terminated with exit code 0.Program failed to restart correctly.I was able to connect with IRGA using Li-Cor software.Powercycled Sonic and program still failed to start correctly.Rebooted computer and program started.RTS at 1430 GMT.

8/18/15

PD

CO2 Flux:Error shown:Program terminated with exit code 0.Program failed to restart correctly.I was able to connect with IRGA using Li-Cor software.Powercycled Sonic and program still failed to start correctly.Rebooted computer and program started.RTS at 1430 GMT.

7/20/15

PAD/JEM

4M CO2 Flux. Mentor is on site to upgrade CO2 flux systems.R/R 4M *IRGA.R/R **Sonic Anemometer.RTS at 2050 GMT.

 

*Old IRGA - SN 75B/H-00157. DOE 6436822.

*New IRGA - Model LI-75H-2591, AIU-1452, DOE 6789966.

**Old Sonic - Model 1352, SN 33, DOE# 6437362.

**New Sonic - Model 1210-PK-085, SN 150601.

7/20/15

PAD/JEM

60M CO2 Flux.Mentor is on site to upgrade CO2 flux systems.Lowered side B booms at 1405 GMT.Removed 60M *IRGA.IRGA is being decommissioned and will be sent to Mentor.R/R **Sonic Anemometer.RTS at 2050 GMT.

 

*Old IRGA - SN 75B/H-0435. DOE 6482157.

** Old Sonic - Model 1352, DOE# 6481280.

**New Sonic -Model 1210-PK-085, SN 150305.

7/20/15

PAD/JEM

25M CO2 Flux.Mentor is on site to upgrade CO2 flux systems.Lowered side B booms at 1405 GMT.Removed *25M IRGA, IRGA is being decommissioned and will be sent to Mentor.R/R **Sonic Anemometer.RTS at 2050 GMT.

 

*Old IRGA Ė SN 75B/H-0469

**Old Sonic - S/N 57, Model 1352, DOE# 6481297

**New Sonic - SN 150602, Model 1210-PK-085,

6/25/15

PD

PGS building upgrade is complete.CO2 Flux computer and network hardware needs to be moved from temporary EF enclosure located base of the 60M tower and moved into the PGS building.Stopped program and shut down computer.Ran conduit to PGS building and ran fibers and power.Moved computer and all network hardware into PGS building and installed.Powered up system, computers and started program.RTS on 6/26/15 at 2005 GMT.

6/10/15

PD

Planned power shutdown to connect transformer located by AOS trailer to distribution panel that was recently installed.Stopped 4m, 25m and 60m Sonic programs at 1306 GMT and shut down computer.Started program after power was restored at 1552 GMT.

6/5/15

PD/MS

Power to all ground mounted transformers is being cut in order to upgrade the transformers at AOS trailer, IDP2 and 50 profiler pad.Instrument needs to be shut down.Stopped 4m, 25m and 60m Sonic programs and shut down CO2 Flux computer.Restarted system after power was restored.RTS at 1834 GMT.

4/16/15

JEM

Notified by Mentor the 25M and 60M IRGA will be removed from service and the 25M and 60M Sonic Anemometer will be left in service.Mentor requests pin out of cables and power/comms box as there is no known wiring diagram available.Lowered side B of tower at 1849 GMT.Developed wiring diagram and took pictures of power/comms box.Raised tower at 1959 GMT.

4/2/15

PD

PGS building is being upgraded.25M and 60M CO2 Flux system was shut down on 4/1/15 at 1738 GMT to move CO2 computer and all related equipment to temporary EF enclosure set at the base of the 60M tower.Systems were powered up on 4/1/2015 but 25M and 60M CO2 Flux did not start up correctly.At 1430 GMT, removed and cleaned the ends of fibers for the 25M and 60M system in the enclosure.Replaced a suspect ethernet cable for the 25M IRGA inside the enclosure and powercycled 25M side B of the tower.Program started correctly.RTS at 1500 GMT.

4/1/15

PD/KT

PGS building is to be replaced with larger building.CO2 Flux and CO2 Met computers need to be moved to temporary EF computer/comms enclosure set at the base of 60M tower.At 1738 GMT, shut down 4, 25, and 60M CO2 Flux computers and comms.Shut down CO2 Met computer and power to 4M systems.Facilities moved fibers from PGS shed to enclosure for 4, 25, and 60M CO2 Flux and 4M CO2 Met systems.Computer department installed UPS and Moxa in enclosure.Ethernet cable was run from enclosure to PGS building for Spectronus.Moved CO2Flux and CO2Met computers, Keyspan serial-USB converters, monitor, keyboard, and mouse switch and 25 and 60M bit drivers from PGS to enclosure.All systems except 25 and 60M CO2 Flux were RTS at 2128 GMT.

3/13/15

JEM

60M IRGA was removed for calibration and needs to be installed.Lowered tower††††††††††††††††††††††††††††††††††††† †††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††

3/13/15

JEM

25M IRGA was removed for calibration and needs to be installed.Lowered tower†††††††††††††††††††††††† ††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††††

3/13/15

JEM

4M IRGA was removed for 6 month calibration and needs to be installed.Installed IRGA and started Sonic 1 program.RTS at 2123 GMT.

3/12/15

JEM

60M IRGA is due for 6 month calibration.Lowered tower and removed IRGA. Raised tower.RTS at 1842 GMT.

3/12/15

JEM

25M IRGA is due for 6 month calibration.Lowered tower and removed IRGA. Raised tower.RTS at 1842 GMT.

3/12/15

JEM

4M IRGA is due for 6 month calibration.Removed IRGA. RTS at 1800 GMT.

3/2/2015

PD/KT

Sonic 1, 2 and 3 programs are not showing graphs.Indicator lights on Keyspan media adapters are on solid for all 3 systems.Rebooted computer.Computer cannot access the network during bootup.Contacted computer department and was informed preferred DNS address had been changed but not on this computer.We do not have root access to this computer so we cannot change DNS address.Emailed Mentor.Waiting on reply.RTS at 1555 GMT.

11/7/14

PD/KT

Farmer has planted wheat in Field 8 South of tower.4M radiometer boom needs to be installed.Moved 4M boom into field at 1445 GMT.Found 2 heat flux plate wires where protective sheath was gone showing exposed wire.Removed HF plates and took to Electronics Lab for repair.Installed repaired HF plates, took soil cores, and installed soil moisture sensors.2 of the soil temperature probes are missing the aluminum probe.Mentor will send replacements.Took soil core samples to shipping dept.RTS at 1845 GMT.

10/28/14

CAM

Lowered the West booms at 1506 GMT for T/H calibration changes.Raised booms and RTS at 2230 GMT.No problem with this instrument.

11/15/13

JEM

During the extensive troubleshooting of the 25M CO2FLUX system yesterday a few wires were broken free from the 25M IRGA bit driver plug.Due to lack of time yesterday the wires were temporarily fixed to allow for data collection but they need to be repaired correctly.Stopped the 25M CO2FLUX data collection @1512 GMT.Made necessary repairs to the bit driver plug.Reconnected and started the 25M CO2FLUX data collection @ 1559 GMT.Status was OK after the restart.RS @1600 GMT.

11/14/13

JEM

The bit drivers that were ordered to fix the 25M and 60M CO2 flux systems have arrived.Replaced the suspect bit drivers but still had communication issues.Lowered the West tower booms at1700 GMT.Connected the umbilical cables to the boom and the power/data systems.Trouble shooting found that the 25M IRGA was not functioning properly.R/R the IRGA.Was able to communicate with the IRGA from the base of the tower, but the Sonic would only stream 1 character ever second or so.The tower remained down most of the day for troubleshooting.Found a wire in the terminal strip junction box on the 25M boom that was not securely crimped in the spade terminal.The wire could be freely pulled out and pushed in to the spade.Re-crimped this connection and was then able to stream data from the Sonic at the base of the tower.Reconnected all fibers/bit drivers and tried to run the data collection software on the computer.The program failed to buffer any data.The activity lights for both the 25M Sonic and 25M IRGA were flashing on the USB/Serial device indicating data was being received.Removed the serial connections from this hub and tried connecting using a laptop and hyper terminal.This revealed that the 25M Sonic and 25M IRGA were both working, but based on cable labels were not connected to the correct spot on the serial hub.Tracing out all cables/fibers/bit drivers revealed that the fibers were mislabeled at the base of the tower.These were original labels that have always been in place.Before all of the troubleshooting and swapping, while the fibers were mislabeled they must have originally been connected to the proper bit driver while being labeled wrong.Removed all labels and relabeled everything properly.There are a couple of wires on one of the bit drivers that broke during all of the troubleshooting.Restarted the collection program and it buffered as expected.The wires were connected in a temporary manner and will be permanently fixed tomorrow due to time today.††† Raised the West booms at 2300 GMT.Started all CO2 Flux data collection programs without issue.Will return tomorrow and fix the wires on the 25 pin IRGA bit driver connector.RTS @2300 GMT.

7/22/13

KT

Farmer was getting ready to work on his field which required the 4 meter flux system to be relocated until the farmer was done. At 1310 GMT relocated the 4 meter flux system to the West guy wire of the 60 meter tower until further notice. RTS at 1330 GMT.

6/26/13

CAM/JEM

The landowner cut and bailed the wheat stubble around the 60m tower.The power/data cable and conduit that runs to the CO2 Met station was snagged by the combine or tractor and ripped from both ends and torn to several pieces.A new cable was built in the lab and pulled through new conduit.Used existing connectors at the instrument end but had to make slight wiring changes (color and pin numbers).Notified mentor of changes so he could update documentation.One of the soil heat flux cables was also severed.No spare sensors available so the cable was spliced.Tested communications and everything appears to be operating as it should.Asked mentor to verify good data and to advise if splicing the heat flux cable is acceptable.RTS at 1640 GMT.

6/17/13

JEM

The 25M and 60M IRGAs have been calibrated and need to be reinstalled.Lowered the West booms @1947 GMT.Reinstalled the IRGAs and tested on the ground before raising the booms.West booms were raised back into place @2135 GMT. Restarted the computer program @2145 GMT.Status OK.RTS @2150 GMT.

6/13/13

JEM

West booms were lowered from 1455 GMT to 1630 GMT so the 25M and 60M IRGAs could be calibrated.IRGAs will be reinstalled at a later date.

2/26/13

PAD

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.Power was restored at 1920 GMT.Sonic 2 and 3 restarted.Powercycled Sonic 1 IRGA and Sonic.Unable to check graphs due to computer will not recognize mouse.Keyspan lights indicate all instruments are communicating with computer.RTS at 2000 GMT.

2/14/13

PAD/KT

Informed by mentor 25M CO2 flux values are out of range.Issues first showed up on 2/6/2013.Mentor requests instruments power cycled to see if that will correct the issue.At 2200 GMT, observed Sonic 2 values for CO2L and H2OL values out of range.Power cycled 25M level, side B of tower.No change in values.Power cycled 25M level again and rebooted CO2 flux computer.No change in values.Troubleshooting at this time.RTS at 2215 GMT.

2/7/13

PAD

CO2flux computer (aka PGS computer) that collects data from 4M met system failed.Mentor sent replacement *computer and computer dept was able to recover and copy all relevant files to replacement computer.Mentor requests computer installed.Assisted Windows administrator installing computer at 1930 GMT.Administrator loaded necessary drivers for Keyspan serial-usb convertors and Picarro shed Nikon camera. Computer will now communicate with camera and 4M met datalogger.Some settings on PC208 for datalogger were not known.Notified Mentor computer was installed and asked that he check datalogger settings.Dell Optiplex 745, Model DCCY, SN F14KRD1, LBLDOE 6658330.RTS at 2100 GMT.

1/31/13

PAD/KT

CO2 MET computer recently has been shutting down and failing to boot up correctly.Computer was removed and taken to lab on 1/30/13 but has failed again.Mentor is sending replacement computer to SGP and requests removed computer be taken to lab to see if files can be recovered from harddrive.Removed *PGS computer and took to computer lab.Dell Optiplex GX260, Model DHM, SN 4LT5221, LBLDOE 6491364.RTS at 2155 GMT.

1/30/13

PAD

CO2 MET computer was found down on 1/29/13.It was restarted at that time but shut down later the same day.Mentor requests computer removed and taken to lab to be checked out and repaired if possible.Removed PGS computer at 1848 and took to computer lab.Windows administrator found corrupt files on C: drive, fixed corrupt files, and was able to boot up computer in lab.Installed computer in PGS shelter but was unable to boot up computer.Notified Mentor.RTS at 2248 GMT.

1/8/12

CAM/JEM

Continued T/S of 4m CO2 Flux T1 sensor/cable problems.Found that the new sensorís T and RH wires were reversed in the connector that was installed by the vendor.Swapped the T and RH pins in the connector and the new sensor now works properly.Also found and repaired multiple wiring and pin failures in both of the original sensorís connectors and wiring harness.†† The failures were due to corrosion that has built up on the connections after many years in the field.After these repairs were made, both of the original sensors are now working and compare well with a calibrated Vaisala T/RH sensor.4m system RTS at 1930 GMT with the original sensors in place and working properly.Mentor will send new connectors, pins and wire so a new wiring harness can be built here in the lab.New T/RH sensor placed into spares storage in PGS shed.

1/7/12

CAM/JEM

4m system T1 sensor is reading 8-10C low and RH is reading ~22% high.Mentor ordered a replacement sensor and had it sent here to be installed.Noted T1 reading 2C and T2 reading 10C at maintenance start time of 1945 GMT (actual air temp is ~10C). R/R T1 sensor but new sensor read -40C.T/S found wiring harness failure.Removed harness to lab and replaced broken pin on T1 connector.Reinstalled harness but now both the old and new sensors read -40C when installed in T1 location.Tried new sensor in T2 location and it read -8C there. Suspect either a problem with new sensor or additional problems with wiring harness.Will continue T/S tomorrow.RTS at 2240 GMT with original sensors.

12/24/12

PAD/JEM

60M CO2 flux only shows data for part of the time.System was not graphing and would not start graphing when command was entered.Rebooted computer.Sonic 1 and 2 restarted correctly but Sonic 3 did not.Connected with Sonic 3 IRGA and Sonic using laptop and Hyperterminal.IRGA scrolls data correctly but Sonic anemometer would only show symbols at very slow rate.R/R bit drivers one at a time with spare for 60M anemometer with no change.Swapped 25M and 60M Sonic bit drivers and fibers.Unable to isolate issue but appears there could be a problem with both 60M anemometer and 60M comms.60M level, side B was powercycled several times during troubleshooting.Troubleshooting at this time.†† RTS at 1836 GMT.

12/18/12

KT

The graph display for the 60M CO2 Flux instrument would not display. Power cycled instrument. When graph display was visible, the values shown on the graph were spiking and appeared to be updating slowly. Power cycled instrument again. Instrument came online and began normal operations. Instrument mentor was notified. RTS at 1608 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.CO2Flux was gracefully shut down at 1322 GMT in anticipation of power outage.Power was restored at 2115 GMT after electrical upgrades were complete.Restarted CO2FLux and RTS at 2142 GMT.

11/12/12

PAD/KT

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.Power was restored at approximately 1930 GMT after all transformers were replaced.At 2110 GMT, rebooted CO2 flux computer and power cycled Keyspan convertors.Sonic 1, 2, and 3 programs all started correctly.RTS at 2130 GMT.

11/1/12

PAD/CM/KT

4M CO2 met radiometer boom was removed from field to allow farmer to plant wheat and needs to be installed back in field.Moved 4M boom back into field.Took soil cores at 10 and 20 cm depth.Installed soil moisture, temperature, and heat flux plates.Installed rain gauge.Verified all soil sensors are showing values.RTS at 1949 GMT.

10/11/12

JEM

The tower needs to be lowered to reinstall the 25M and 60M IRGA's.Lowered the booms @1753 GMT.Installed and tested the IRGAís.The 60m IRGA power cable has failed.Made necessary repairs to the power cable and tested OK.The booms were raised and RTS @1915 GMT.

10/10/12

CAM

D. Cook

West booms needed to be lowered for 60M T/H maintenance.Lowered the west booms at 1905 GMT.Raised booms and RTS at 2045 GMT.No problem with this instrument.

10/9/12

CAM

D. Cook

Booms needed to be lowered for for annual 25/60m T/H calibration changes.Lowered the west booms at 1845 GMT.Removed 25 and 60m IRGAs for calibration.IRGAs should be reinstalled on 10/11. Raised booms and RTS at 2210 GMT.

10/3/12

KT

The 4m CO2 Flux system needed to be moved so the land owner can work his field.The sensors were dug up and tie wrapped to system frame, the rain gauge was disconnected and placed inside the PGS building. Moved the 4m CO2 Flux system.

8/16/12

PAD

4M IRGA was removed for a scheduled calibration 8/15/12.Calibration was completed and sensor needs to be installed.Installed IRGA at 1800 GMT.After install was unable to connect to IRGA.Troubleshooting found insulation had deteriorated behind knurled nut on power cord at end that connects to sensor causing broken wire.Removed cord and took to electronics lab for repair.Installed repaired power cord and was able to connect with 4M system.SN 75 B/H-0157.RTS at 1920 GMT.

8/15/12

PAD

4M IRGA is due for a scheduled calibration.Stopped Sonic 1 program at 1606 GMT and removed IRGA for calibration.SN 75 B/H-0157.RTS at 1626 GMT.

6/19/12

PAD

4M radiometer boom and all soil sensors were removed to allow farmer to harvest wheat.Farmer then planted cow peas in wheat stubble.4M boom needs to be moved back into field and sensors installed.Moved 4M boom back into field at 1340 GMT.Installed soil moisture, temperature, and heat flux plates.Took soil cores.Installed rain gauge tipping bucket and connected all sensors.RTS at 1540 GMT.

5/22/12

PAD

Farmer has started harvesting wheat in field where 60m tower is located.4m radiometer boom needs to be removed.Disconnected all soil sensors and rain gauge from 4m boom.Removed all sensors and boom from field.RTS at 1330 GMT.

3/8/12

PAD

Radon Gas Analyzer inlet pipe has separated approximately 40M above the ground.Contractor is on site today to climb tower and repair inlet tube.Power to tower needs to be shut down before contractor can climb. Stopped Sonic 2 program and shut down power at base of tower to 25 and 60M levels, side B at 2023 GMT.After repairs were made and complete inspection of inlet pipe was performed, power was restored and Sonic 2 program restarted.No issues with this instrument are known at this time.RTS at 2257 GMT.

2/14/12

PAD/CAM

4M IRGA was removed on 2/7/12 for calibration and needs to be installed.Installed IRGA on 4M tower and started Sonic 1 program.SN 75 B/H-0157.RTS at 2125 GMT.

2/14/12

PAD/CAM

60M IRGA was removed on 2/7/12 for calibration and needs to be installed.Lowered tower at 2023 GMT and installed IRGA.Raised tower at 2105 and started Sonic 3 program.SN 75 B/H-0128.RTS at 2108 GMT.

2/14/12

PAD/CAM

25M IRGA was removed on 2/7/12 for calibration and needs to be installed.Lowered tower at 2023 GMT and installed IRGA.Raised tower at 2105 and started Sonic 2 program.SN 75 B/H-0435.RTS at 2108 GMT.

2/7/12

PAD

Mentor requests 4M IRGA calibrated.Stopped Sonic 1 program at 1504 GMT.Removed IRGA SN 75 B/H-0157.Took IRGA to lab for calibration.RTS at 1524 GMT.

2/7/12

PAD

Mentor requests 25M and 60M IRGAs calibrated.Stopped Sonic 2 and 3 programs and lowered side B booms at 1414 GMT.Removed 25M IRGA SN 75 B/H-0435 and 60M IRGA SN 75 B/H-0128.Raised tower at 1500 GMT.Took IRGAs to lab for calibration.RTS at 1514 GMT.

12/29/11

PAD

Computer updates were performed on CO2 met computer yesterday.At 1635 GMT, notice was shown on monitor stating:Computer updates are complete.Reboot now, Reboot later.†† CO2met values were not believable.Looking at yesterday's data, data stopped around 1900 GMT.Rebooted computer and connected with datalogger.CO2met values now seem to be in reason.RTS at 1650 GMT.

12/22/11

PAD

While checking CO2 flux graphs, computer froze.Graph 2 was still displayed and scrolling, mouse would move but nothing responded.Power cycled computer.After startup, all Sonic programs are now running.RTS at 2010 GMT.

11/14/11

PAD

Field 8 has been planted to wheat.4m boom was moved back into field but was unable to bury soil sensors as it was raining.Soil moisture and soil temperature sensors need to be installed.Dug trenches and took soil cores at 10 and 20 cm depth.Installed soil moisture sensors at those levels.Installed soil temperature sensors at 5, 15, and 25 cm depth.RTS at 2130 GMT.

11/7/11

PAD

Field has been planted to wheat and 4m radiometer boom needs to be moved back into field.Moved 4m boom into field with assistance and installed HF plates and rain gauge.Raining at this time so soil moisture and soil temperature probes were not installed.RTS at 1450 GMT.

11/1/11

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration change outs.†† Maintenance will also affect the 25m and 60m CO2 flux system.Lowered the West booms at 1758 GMT for T/H calibration changes.Raised booms and RTS at 2142 GMT.No problem with this instrument.

10/17/11

PAD

Farmer had tilled part of field 4m radiometer boom is located in and worked around boom.Some parts of the field are still too wet to till.Boom needs to be removed so farmer can till where boom was located when he returns.Removed HF plates and rain gauge.Moved boom out of field.RTS at 1800 GMT.

9/14/11

PAD

System was removed on 9/1/11 to allow farmer to till field and needs to be installed.Moved 4m radiometer boom back into field and installed soil heat flux plates and connected rain gauge.Did not install soil temperature or soil moisture sensors as farmer intends to till field again soon.RTS at 2015 GMT.

9/1/11

PAD

Farmer intends to till field where 4m radiometer boom is located.Disconnected all soil sensors and moved boom out of field.RTS at 1951 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.Power was restored on 8/9/11 at 0211 GMT.Computer has power but was unable to connect with the network which prevents it from starting Sonic programs.Powercycled computer and it connected with network.Sonic 1,2 and 3 programs started correctly.RTS at 1430 GMT.

7/6/11

PAD

4m radiometer boom was removed on 6/29/11 to allow farmer to till field.System needs to be moved back into field.Moved 4m radiometer boom back into field.Took soil cores and installed rain gauge, heat flux plates, soil moisture, and soil temperature sensors.RTS at 1600 GMT.

6/29/11

PAD

Farmer is going to till field 4m radiometer boom is located in.Boom needs to be removed.Removed 4m boom and all soil sensors.RTS at 1350 GMT.

5/26/11

PAD

Notified by mentor 4m CO2 met TGRG has not recorded any precipitation for approximately 6 weeks.TBRG was removed on 3/24/11 and cable was replaced.At 1820 GMT, found signal wires would complete a circuit.Removed funnel and found connectors on ends of signal wires were making contact.Removed one connector and trimmed off outside of ring so connectors would no longer touch.Installed TBRG and tipped 3 times at 1842 GMT.RTS at 1845 GMT.

5/17/11

PAD

Animals had previously damaged 4m system #2 and 4 soil moisture sensors.Sensors were removed, repaired, and placed into service.Sensors need to be removed to install flexible conduit to prevent future damage.Disconnected and removed sensors 2 and 4.Took sensors to lab to remove connector to allow conduit to be installed.RTS at 1741 GMT.

5/11/11

PAD

Soil moisture sensors 1 and 3 were repaired and installed on 5/10/11.Something is causing sensor 3 to drop voltage at datalogger.Troubleshooting found a short in the cable for sensor 3.Sensor was removed and taken to lab where cable was replaced.†† RTS at 1721 GMT.

5/10/11

PAD

Rodents have damaged soil moisture sensor cables 1 and 3 on 4M system.Sensors and cables were removed, repaired, put in protective conduit, and need to be installed.Installed sensors 1 and 3 at 20cm and 10cm respectfully.Connected sensors to system.Unable to connect with datalogger.Troubleshooting found a problem with sensor 3 causing voltage drop to datalogger.Left sensor 3 disconnected.Troubleshooting at this time.RTS at 2152 GMT.

4/26/11

PAD

60M IRGA was removed on 4/26/11, calibrated, and needs to be installed back on 60M tower.Lowered tower at 1721 GMT.Installed IRGA, SN 75 B/H-0128.Raised tower at 1800 GMT and started Sonic3 program.RTS at 1821 GMT.

4/26/11

PAD

25M IRGA was removed on 4/26/11, calibrated, and needs to be installed back on 60M tower.Lowered tower at 1721 GMT.Installed IRGA, SN 75 B/H-0435.Raised tower at 1800 GMT and started Sonic 2 program.RTS at 1821 GMT.

4/26/11

PAD

Mentor requests 60m IRGA calibrated.Stopped Sonic 3 program and lowered side B tower.Removed IRGA and took to GIF trailer for calibration.SN 75 B/H-0128.RTS at 1434 GMT.

4/26/11

PAD

Mentor requests 25m IRGA calibrated.Stopped Sonic 2 program and lowered side B tower.Removed IRGA and took to GIF trailer for calibration.SN 75- B/H-0435.RTS at 1434 GMT.

4/20/11

PAD

IRGA was removed for calibration on 4/19/11 and needs to be installed and RTS.Installed IRGA, SN B/H-0157 at 1844 GMT.Connected all cables and started Sonic 1 program.RTS at 1904 GMT.

4/19/11

PAD

Animals have cut #1 and #3 soil moisture sensor cables.Removed cables and sensors and took pictures of damage to send to mentor.Working on plan with mentor to put cables in conduit for protection at this time.RTS at 1845 GMT.

4/19/11

PAD

Mentor request 4m IRGA calibrated.Stopped Sonic 1 program and removed IRGA, SN 75 B/H-0157.Took sensor to GIF trailer to start calibration.RTS at 1339 GMT.

3/24/11

PAD

On 3/23/11, observed rodent damage to 4m met system rain gauge cable.Removed cable and took to lab for repair.Cable was replaced and installed rain gauge.RTS at 1950 GMT.

3/24/11

PAD

4m radiometer boom was removed from field on 3/23/11 to allow farmer to apply herbicide and needs to be moved back into field.Moved radiometer boom back into field and connected all sensors.Buried HF plates and verified all sensors are producing data.RTS at 1950 GMT.

3/23/11

PAD

Farmer intends to spray field 4m radiometer boom is located.Disconnected sensors and moved 4m boom out of field at 1250 GMT.

2/16/11

PAD

Notified by mentor soil moisture and other values from 4m met system were not correct.At 1904 GMT, found rodent had cut wires on sensors 2, 3, and 4.Removed all soil moisture sensors and took to lab.Repaired all cables and reinstalled in field, 1 and 3 at 20 cm, 2 and 4 at 10 cm.RTS at 2204 GMT.

10/19/10

CAM

Tower Systems on site for tower inspections.Need to secure power for maintenance.Secured AC power to 25m and 60m levels at 1407 GMT.Restored power following inspection at 1620 GMT.

10/18/10

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration change outs.†† Maintenance will also affect the 25m and 60m CO2 flux system.Lowered the West booms at 1423 GMT for T/H calibration changes.Raised booms and RTS at 2026 GMT.No problem with this instrument.

9/29/10

PAD/CAM/JEM

System was removed on 9/13/10 to allow NH3 fertilizer to be applied and farmer to plant crop.4m boom needs to be installed.At 1920 GMT, moved 4m boom back into field and installed all soil sensors.Soil cores were taken to be sent to LBNL. Verified all sensors are showing values.RTS at 2030 GMT.

9/13/10

PAD

Farmer intends to apply NH3 fertilizer to field.Removed 4M radiometer boom at 1720 GMT.

9/7/10

PAD

4m radiometer boom was removed on 9/3/10 to allow farmer to disk field.Moved radiometer boom back into field at 1730 GMT and installed HF plates.RTS at 1750 GMT.

8/11/10

PAD

*#1 HF plate on 4m radiometer boom is showing -9999s.Troubleshooting found sensor had failed.Installed new HF plate at 1420 GMT.Sensor is still showing -9999s and #2 HF plate is now sporadically showing -9999s.Disconnected sensor wires at CR10 wiring panel and tested sensors with voltmeter.Sensor responded correctly depending on what side of sensor was facing the sun.Reconnected sensors to wiring panel.Both sensors are now showing believable values.Sensors will not be installed in ground until crop is planted in field where boom is located.Notified instrument mentor so sensor sensitivity could be changed in datalogger program.RTS at 1610 GMT.

*New sensor type: HFP01.SN 003854.Sensitivity 65.1 uV/Wm2.

8/10/10

PAD

4M radiometer boom was removed on 8/4/10 to allow farmer to burn field and needs to be moved back into field.Move 4M boom back into field at 1325 GMT.Did not reconnect soil sensors as farmer is planning to till field after as soon as we receive enough rain.RTS at 1345 GMT.

8/5/10

PAD

4M sonic and IRGA were removed to allow farmer to burn wheat stubble on 8/4/10.Instruments need to be put back into service.Installed sonic, LBLDOE 6437362, and IRGA, SN 75H-0157 at 1401 GMT.Started Sonic 1 program.RTS at 1416 GMT.

8/4/10

PAD

Farmer intends to burn field 4M CO2flux system is located in.Mentor requests 4M sonic and IRGA removed.Stopped Sonic 1 program at 1836 GMT and removed Sonic, LBLDOE 6437362, and IRGA, SN 75H-0157.RTS at 1851 GMT.

8/3/10

PAD

Farmer is going to burn wheat stubble in field 4m system is located.Removed 4m radiometer boom and all soil sensors at 1345 GMT.RTs at 1415 GMT.

6/25/10

PAD

Wheat field that 4m radiometer boom is located in was harvested on 6/22/10.Radiometer boom was left in place and flagged.Wheat standing under boom needs to be harvested by hand to mimic surrounding field that was harvested with a stripper header.At 1400 GMT, assisted vegetation technician in removing wheat heads under boom, base of tower, and area surrounding tower anchors.

6/11/10

PAD

60M tower cameras are not working.Computer department requests both sides of tower power cycled at the 60M level to reset cameras.Power cycled Side A and Side B at 60M level at 1738 GMT.Left Sonic 3 program running and data is still graphing after power cycle.No problem with this instrument known at this time.RTS at 1753 GMT.

5/17/10

PAD

CF lost power on 5/16/10 between 0952-1504 GMT.At 1256 GMT, observed message on CO2 flux computer:Could not look up internet address for local host. local domain.This will prevent GNOME from operating correctly.Try again or Cancel.Mouse is not responsive.Internet service at CF was later restored.At 1546 GMT, rebooted CO2 flux computer.Sonic 1,2, and 3 programs all started correctly.RTS at 1601 GMT.

5/11/10

PAD

4M radiometer boom was removed on 5/6/10 to allow farmer to apply fungicide to wheat field and needs to be installed.Moved 4M boom back into field, buried heat flux plates and connected all soil sensors at 1800 GMT.RTS at 1820 GMT.

5/6/10

PAD

Farmer intends to spray wheat with fungicide.4m radiometer boom needs to be removed from field to allow spraying.Disconnected soil sensors from boom and removed at 1911 GMT.RTS at 1931 GMT.

4/8/10

PAD

4m radiometer boom was removed from field to allow farmer to apply fertilizer and herbicide to wheat earlier today and needs to be placed back into field.Moved 4m boom back into field and connected all soil sensors.RTS at 1835 GMT.

4/8/10

PAD

4m radiometer boom needs to be removed from field to allow farmer to apply fertilizer and herbicide to wheat.Disconnected soil sensors and removed 4m boom.RTS at 1250 GMT.

4/7/10

PAD

4m IRGA was removed on 3/30/10 and sent to LiCor for repairs and calibration.Instrument needs to be installed.Installed *IRGA and connected laptop to instrument.Changed the following settings on RS232 tab:Baud rate 9600 to 38400.Frequency 1 to 10.All data boxes were selected.Selected Temperature, Pressure, CO2 (mmol/m3), H20(mmol/m3), and Diagnostics.Unchecked options and selected both (ODOA).On output tab:Changed delay from 0 to 11.DA#1 none changed to H20(mmol/m3).DA#2 changed none to CO2(mmol/m3).All other settings are correct.Connected IRGA to CO2flux computer and started Sonic 1 program.Program started correctly.Notified mentor.RTS at 1440 GMT.*New- 75 B/H-0157

3/30/10

PAD

4m IRGA readings for CO2 and water vapor are out of range.Mentor requests *4m IRGA removed and sent to LiCor for repairs.Removed IRGA and took to shipping dept. with RMA provided by mentor.RTS at 1609 GMT.*Old - B/H- 0157.†† Model LI-7500

1/27/10

PAD

4m boom was removed on 1/26/10 to allow farmer to apply fertilizer in wheat field instrument is located in.Boom needs to be placed back into service.Move 4m boom back into field and reconnected all soil sensors.RTS at 1930 GMT.

1/26/10

PAD

Notified by farmer he intends to apply fertilizer in the field 4m boom is located.†† Disconnected soil sensors and removed 4m boom.RTS at 2125 GMT.

12/16/09

PAD

#2 HF plate needs to be replaced with spare provided by mentor.Installed HF #2, Type HFP01, SN 003856.Notified mentor of installation so calibration coefficient can be changed in program to reflect new HF plate.RTS at 2132 GMT.

11/18/09

PAD

#2 HF plate is showing -9999 values.Previous troubleshooting indicated CR10X wiring panel is faulty and needs replaced.Shut off power to system, removed datalogger box, and took to lab.Lab technician R/R wiring panel.Installed datalogger box and powered up system.There was no change to #2 HF plate value.Further troubleshooting found bad HF plate.Waiting on replacement at this time.RTS at 2113 GMT.

10/27/09

PAD

4m system HF #2 values are either not in range or shows as -9999.Cleaned connection at port box, reseated HF2 wires at wiring panel, physically moved wires at wiring panel and connector to try and get HF2 number to fluctuate with mixed results.Removed data logger and reseated with no change.Wiring panel looks like it has arced between two solder joints on the back side.When the connector bar for the HF wires at wiring panel is pressed, number fluctuates between -9999 and a number between -8 and -60.Suspect wiring panel needs to be replaced.Notified mentor.RTS at 1900 GMT.

10/20/09

PAD

Heat flux plates were installed on 4m system on 10/19/09.#2 is alternating between -9999 and a number usually lower than other three HF plates.Suspect corrosion on common connector for HF plates is causing bad readings for #2 HF plate.Removed connector and cleaned pins with small screwdriver and contact cleaner.No change to #2 HF plate readings.Troubleshooting at this time.RTS at 1914 GMT.

10/19/09

PAD

4m CO2 flux heat flux plates need to be installed.#2 and #3 HF plates have been repaired.At 1820 GMT, connected HF plates to system.#4 HF plate shows -9999.Reseated connector and #4 is now registering value in range but #2 is now registering -9999.Checked #2 HF at connector with voltmeter and it appears to be working.Installed HF plates at 5 cm depth.#2 is showing values but is not in line with other 3 HF plates and is showing -9999 most of the time.RTS at 1920 GMT.

10/6/09

PAD

Heat flux plates need to be installed on 4m CO2 flux system.†† Put new conduit on plate wires and installed on system.#2 and #3 plates are registering -9999. Plates will have to be removed to find problem with connector or previous repair splice.RTS at 2120 GMT.

9/30/09

PAD

4m boom was removed on 9/25/09 to allow farmer to fertilize and plant field and needs to be placed back into field.Moved 4m boom into field at 1810 GMT.Took soil cores and installed soil moisture and soil temperature probes.RTS at 1945 GMT.

9/21/09

CAM

Lowered the West booms at 1506 GMT for T/H calibration changes.Raised booms and RTS at 1750 GMT.No problem with this instrument.

8/26/09

PAD

Farmer has finished disking field.8 and 4m boom needs to be redeployed.Moved 4m radiometer boom back into field, aligned E-W and checked level of radiometers.RTS at 1430 GMT.

8/25/09

MK

Farmer intends to disk field.4m boom needs to be removed.Removed boom at 1400 GMT.

8/4/09

PAD

Computer department received spare hard drive for CO2 flux computer from mentor.Mentor requests files and operating system on current hard drive copied to spare.Stopped sonic 1, 2, and 3 programs and shut down computer.Computer department copied hard drive files and operating system onto spare.†† After transfer was complete, spare hard drive was checked to make sure computer would boot up correctly.No problems found.Booted up computer with original drive and Sonic 1, 2, and 3 started correctly.RTS at 1928 GMT.

7/15/09

PAD

4m radiometer boom was removed from field on 7/10/09 to allow farmer to disk.Boom needs to be RTS.Moved 4m boom back into field at 1340 GMT.Did not bury soil sensors since farmer is now tilling field.Soil sensors will be installed after next crop is planted.RTS at 1400 GMT.

7/13/09

PAD

Computer hard drive was R/R on 7/7/09.Mentor cannot remotely access computer.Computer department set IP address and rebooted computer.Computer will now access internet.Notified mentor who verified he can remotely access computer.RTS at 1850 GMT.

7/10/09

PAD

Notified by farmer he intends to disk field 4m radiometer boom is located in.Boom and all soil sensors need to be removed from field.At 1415 GMT, removed all soil sensors, rain gauge, boom, and boom anchors.RTS at 1445 GMT.

7/7/09

PAD

After installing 25 and 60M IRGAs on 7/6/09, computer would not open a new terminal to start Sonic 2 and 3 programs.Rebooted computer and it failed to start up correctly.Several error messages are shown including:Screen Init failed because of version mismatch.[dri].128.0.Kernel module version is 2.1.6 but 2.2 or greater is needed.Fatal I/O error 104.This affects 4, 25, and 60M CO2 Flux systems.At 1525 GMT, computer department troubleshot and found hard drive failure.Spare hard drive in computer provided by mentor was put into service.Computer booted up and all Sonic programs started correctly.RTS at 1615 GMT.

7/6/09

PAD

4M IRGA, SN 78B/H-0157 has been calibrated and needs to be RTS.Installed IRGA and started Sonic 1 program.RTS at 2008 GMT.

7/6/09

PAD

Mentor request 4M IRGA, SN 78B/H-0157 calibrated.Stopped Sonic 1 program and removed IRGA.RTS at 1949 GMT.

7/6/09

PAD

60M IRGA was removed on 7/1/09 for calibration and needs to be RTS.Lowered tower side B at 1848 GMT.Installed 60M IRGA, SN 00128, and control box.Raised tower and tried to start Sonic 3 program but computer would not open new terminal and failed to reboot properly.Tower was RTS at 1938 GMT but Sonic 3 is not working and troubleshooting continues at this time.

7/6/09

PAD

25M IRGA was removed on 7/1/09 for calibration and needs to be RTS.Lowered tower side B at 1848 GMT.Installed 25M IRGA, SN 0435, and control box.Raised tower and tried to start Sonic 2 program but computer would not open new terminal and failed to reboot properly.Tower was RTS at 1938 GMT but Sonic 2 is not working and troubleshooting continues at this time.

7/1/09

PAD

Mentor requested 60M IRGA calibrated.Stopped Sonic 3 program at 1424 GMT, lowered side B booms and removed 60M IRGA, SN 75B/H-0128.RTS at 1604 GMT.

7/1/09

PAD

Mentor requested 25M IRGA calibrated.Stopped Sonic 2 program at 1424 GMT, lowered side B booms and removed 25M IRGA, SN 75B/H-0435.RTS at 1604 GMT.

6/24/09

PAD

4m CO2 flux radiometer boom was removed on 6/18/09 to allow farmer to harvest surrounding wheat crop.Boom needs to be placed back into field and returned to service.At 1527 GMT, used a weedeater to remove some of the crop to try and match surrounding area after farmer harvested the wheat.Moved 4m boom back into place and connected all soil sensors.Checked level of radiometers.RTS at 1627 GMT.

6/18/09

PAD

Notified by farmer they are planning on harvesting field 4m radiometer boom is in.Boom needs to be removed.Disconnected soil sensors, removed 4m boom, and flagged soil probes at 1717 GMT.RTS at 1737 GMT.

3/18/09

PAD

4m radiometer boom was removed on 3/16/09 to allow farmer to apply fertilizer to field.System needs to be moved back into field.At 1747 GMT, we moved 4m boom back into place and connected all soil sensors.Verified all sensors are producing data.RTS at 1807 GMT.

3/16/09

PAD

Notified by farmer he intends to spray field that 4m radiometer boom is located in with pesticide, herbicide, followed by fertilizer at a later time.At 2000 GMT, we disconnected all soil sensors, removed 4m boom, and flagged soil sensors with caution tape.RTS at 2020 GMT.

3/13/09

PAD

Mentor requests *par lite sensor that was temporarily installed on 4m radiometer boom removed.At 1947 GMT, removed Kipp and Zonen sensor RTS at 2002 GMT.*Old - 060850

1/22/09

PAD

Mentor requests uplooking PAR sensor removed for calibration and replaced with sensor from P2 system.He also requests Par lite sensor installed.At 1554 GMT, removed uplooking Par sensor.SN is faded and unreadable.Installed uplooking Par sensor from P2 sensor.SN is faded and unreadable.Installed Par lite sensor.Contacted mentor at 1654 GMT, who remotely loaded updated data logger program with new cal coefficientsfor uplooking Par andPar lite sensors.Connected sensors to data logger at 1703 GMT.†† Checked sensor values on monitor computer.RTS at 1710 GMT.

1/2/09

PAD

CO2 Flux computer will not display on monitor.Hard drive indicator light is on solid.Computer is not responsive to keyboard or mouse.Power cycled computer and it rebooted correctly.Sonic 1, 2, and 3 programs all restarted correctly.RTS at 1649 GMT.

12/22/08

PAD

West side tower booms need to be lowered for 25M T/H maintenance.Stopped all Sonic programs at 1955 GMT.Restarted programs at 2035 GMT.No problems with this instrument.

12/15/08

PAD

West side tower booms need to be lowered to T/S problem with 25M T/H probe.Stopped Sonic 3 program at 1515 GMT.Restarted program at 1627 GMT.No problems known with this instrument.RTS at 1630 GMT.

11/18/08

PAD

Facilities department needs to lower West side 25 and 60M booms to install camera on tower.Stopped Sonic 2 and 3 programs at 1528 GMT.Program was restarted at 1806 GMT.

11/13/08

PAD

Facilities department needs to lower W side booms to mount tower camera.Stopped Sonic 2 and 3 at 1445 GMT.Lowered tower at 1450 GMT.Raised tower at 1537 GMT.No problems with this instrument known.Restarted Sonic 2 program.RTS at 1540 GMT.

11/6/08

PAD

Mentor requested temporary par lite sensor removed from 4m boom.Removed par lite sensor at 1626 GMT.Switched upwelling and downwelling par sensors back to original positions.RTS at 1646 GMT.

10/27/08

PAD

Port #3 in fiber mux located in either PGS shed or base of instruments has failed.Mentor requests communication issue troubleshot and failed mux replaced.T/S found #3 port on fiber mux located at base of instruments is bad.R/R with fiber mux provided by mentor.Removed mux will be sent to mentor for repairs.RTS at 1938 GMT.

9/29/08

PAD

4m CO2 flux radiometer booms and soil sensors were removed on 9/25/08 to allow farmer to harvest corn and plant wheat into corn stubble.At 1445 GMT, technicians moved 4m boom into field.We collected soil cores and installed all soil sensors.RTS at 1645 GMT.

9/25/08

CAM

Need to move 4m CO2 Flux system so farmer can harvest corn and plant wheat.Removed power at 1503 GMT.Dug up soil sensors and moved entire system out of field.

9/19/08

PAD

On 9/12/08, bit driver from 25m IRGA was removed for use in 60m system as per mentor's request.Replacement bit driver was ordered and needs to be installed.Installed bit driver for 25m IRGA and started Sonic 2.Program started correctly and all values look good.Notified instrument mentor.RTS at 2042 GMT.

9/9/08

PAD

Sonic 3 program will not display graph for 60m data.This has been an ongoing issue.After restarting program, it will usually stop after about 1 hour.Media converter box indicates communication between computer and 60m system.Stopped and exited Sonic 3.At 1517 GMT, power cycled instrument from base of tower.†† Unable to restart Sonic 3, Segmentation Fault error is given.At 1740 GMT, I was unable to connect with IRGA using LiCor software but was able to connect with Hyperterminal.Sent (Program Reset (True)) message to IRGA.Instrument will now communicate with laptop and was able to connect using LiCor software.Monitored values looked good.Checked communication with Sonic Anemometer using Hyperterminal and laptop.Reconnected IRGA and Sonic to CO2 Flux computer at media converter.Sonic 3 program restarted correctly and will now graph data.Will continue to monitor.Notified instrument mentor.RTS at 1810 GMT.

9/8/08

PAD

Notified by mentor 60m CO2 Flux system is not collecting data.Media converter box indicates instrument is communicating with computer.Sonic 3 program will not display graph of data.Terminal shows:Graph Display off!./ get Sonic 3:Line 3:4271 Segmentation fault.Exited/ restarted Sonic3program.Graphs will now display on monitor computer.RTS at 2002 GMT.

9/5/08

PAD

Last data collected for 4m CO2 met system was on 8/26/08. Troubleshooting found internal problem with 4m CO2 met datalogger.Removed CR10X, SN 21686, and installed SN 32176.Loaded DLD program and RTS at 1600 GMT.

9/2/08

MK

Last data collection from 4m CO2 met data system was on 8/26/08.Data logger has 9.2V at input and fluctuating around 11.5V when disconnected.Troubleshooting found water in conduit between power supply box for 4m system and instrument tower for 4m IRGA and Sonic.Removed conduit and found nick in power wire insulation.Repaired power wire and reinstalled conduit to prevent water entry.Datalogger now shows 11.5V.RTS at 1945 GMT.

9/4/08

PAD

Sonic 3 is not graphing.Media converter box indicates computer is communicating with 60M IRGA and Sonic but no graphs are shown.Exitedand restarted program.Sonic 3 graph is now updating.RTS at 1558 GMT.

8/28/08

PAD

While doing PM rounds, computer would connect with 4m CO2met datalogger but values will not update.Datalogger battery voltage shows 9.4V.Checked voltage at power supply and it has a steady 11.9V output.†† Voltage at datalogger input power measures 10.0V.When power wire is unplugged from datalogger, it measures 11.9V.Reconnected power wire to datalogger and it shows 10.0-10.4V at input connector.Monitor shows 10.0V and it measures 11.9V at power supply.Troubleshooting at this time.†† RTS at 2100 GMT.

8/26/08

PAD

Sonic 3 is not graphing.Status shows:Line 328276 Segmentation fault.Exited/ restarted Sonic 3 program.Sonic 3 started correctly and will now graph.RTS at 1555 GMT.

8/26/08

PAD

4m Sonic 1 program is not graphing data.Media converter indicates computer is communicating with instrument but no graph is shown.Last status shows :Line31487 Segmentation fault.Exited/ restarted program.Program was very slow to buffer and start but would still not show graphs.Power cycled Sonic and IRGA and restarted Sonic 1.Status shows:†† Line 37593 Segmentation fault.Power cycled Keystone media converter for 4m system and restarted Sonic 1.Sonic 1 started correctly and will now graph.RTS at 1605 GMT.

7/31/08

PAD

4m CO2 Flux radiometer boom was removed from field to allow farmer to spray corn field with herbicide and needs to be installed.Four technicians moved boom into field and connected all soil sensors.RTS at 1603 GMT.

7/23/08

PAD

Notified by farmer they will spray field 4m system is located in with Roundup herbicide tomorrow.4M CO2 Flux radiometer boom needs to be removed from field.Three techs and I disconnected all soil sensors, power and comm cables from boom and removed from field. RTS at 1420 GMT.

7/15/08

PAD

CF lost power on 7/13/08 between 0028 and 0137 GMT.Instrument lost power but computer did not shut down.Notified by mentor on 7/15/08 that instrument was not operating.†† Exited and restarted Sonic 3.Program restarted correctly.RTS at 1450 GMT.

7/11/08

PAD

4m CO2 radiometer boom was removed on 7/8/08 so the farmer could apply fertilizer needs to be moved into field.Moved 4m boom out into field and connected all soil sensors.RTS at 1907 GMT.

7/8/08

PAD

4m CO2 radiometer boom needs to be removed from field so the farmer can apply fertilizer.Disconnected all soil sensors and moved 4m boom out of field.RTS at 1430 GMT.

6/30/08

PAD

Tower Systems was on site on 6/28/08 for tower maintenance.Sonic 2 and Sonic 3 programs were stopped at 1203 GMT and power to tower was shut down.Restarted Sonic 2 and Sonic 3 programs on 6/30/08 at 1447 GMT.RTS at 1502 GMT.

6/18/08

PAD

CF experienced power outages on 6/16/08.Power was restored at 1947 GMT.Notified by mentor 60M Flux system did not restart correctly.††† Exited Sonic 3 program and restarted.Program started correctly and is graphing data.Notified instrument mentor.RTS at 1505 GMT.

6/2/08

PAD

CF experienced extended power outage on 6/1/08.Reset Keyspan converters and restarted CO2 Flux computer.All Sonic programs restarted.Notified instrument mentor.RTS at 1330 GMT.

5/29/08

PAD

4m radiometer boom had been removed from field on 5/22/08 to allow farmer to apply fertilizer.System needs to be moved back into place and sensors connected.At 1730 GMT, we moved boom back into field and connected all soil sensors and rain gauge.RTS at 1750 GMT.

5/29/08

PAD

Mentor requests 60m IRGA calibrated.Mark Klassen and Ifinished calibration of IRGA, SN 75 B/H-0128.Lowered side B, West side booms at 1926 GMT and installed IRGA.Raised booms at 2006 GMT and started Sonic 3 program.RTS at 2006 GMT.

5/29/08

PAD

Mentor requests 25m IRGA calibrated.Mark Klassen and I finished calibration of IRGA, SN 75 B/H-0435.Lowered side B, West side booms at 1926 GMT and installed IRGA.Raised booms at 2006 GMT and started Sonic 2 program.RTS at 2006 GMT.

5/29/08

PAD

Mentor requests 4m IRGA calibrated.†† Mark Klassen and Ifinished calibration of IRGA, SN 75 B/H-0157.Installed instrument on 4m system and started Sonic 1 program.RTS at 1920 GMT.

5/28/08

PAD

Mentor requests 60m IRGA calibrated.Stopped Sonic 2 program and lowered side B, West side 25m boom.Removed IRGA, SN 75 B/H-0128.†† Mark Klassen and I did calibration procedures that can be done today and will finish tomorrow after IRGA runs 24 hrs.

5/28/08

PAD

Mentor requests 25m IRGA calibrated.Stopped Sonic 2 program and lowered side B, West side 25m boom.Removed IRGA.†† Mark Klassen and I did calibration procedures that can be done today and will finish tomorrow after IRGA runs 24 hrs.

5/28/08

PAD

Mentor requests 4m IRGA calibrated.Stopped Sonic 1 program and removed IRGA.†† Mark Klassen and I did calibration procedures that can be done today and will finish tomorrow after IRGA runs 24 hrs.

5/22/08

PAD

Farmer needs to apply fertilizer to field 8. 4m CO2 flux radiometer boom needs to be removed.At 1400 GMT, labeled instrument wires and unplugged from system.Removed 4m boom and flagged soil sensors.RTS at 1445 GMT.

5/19/08

PAD

4m CO2 flux system was removed to allow farmer to plant field.System needs to be moved back into the field.At 1400 GMT, we moved the 4m radiometer boom back into field.Took soil cores and installed soil moisture probes, soil temperature probes, and heat flux plates.RTS at 1630 GMT.

4/29/08

PAD

CF experienced extended power outage on 4/26/08 at 0034 GMT.Power was restored at 0459 GMT.At 1538 GMT, found computer has power but is not running.Started computer but Sonic programs failed to start correctly.Power cycled serial/ USB converters, and instruments on 25 and 60m level.††† Sonic 1 data was streaming on Sonic 2 port on converter.Shut down computer and power cycled instruments.All Sonic programs restarted correctly.RTS at 1608 GMT.

4/29/08

PAD

CF experienced extended power outage on 4/26/08 at 0034 GMT.Power was restored at 0459 GMT.At 1538 GMT, found computer has power but is not running.Started computer but Sonic programs failed to start correctly.Power cycled serial/ USB converters, and instruments on 25 and 60m level.†† Sonic program still failed to start.Sonic 1 data was streaming on Sonic 2 port on converter.Shut down computer and power cycled instruments.All Sonic programs restarted correctly.RTS at 1608 GMT.

4/29/08

PAD

At 1538 GMT, found computer has power but is not running.Started computer but Sonic program failed to start correctly.Started Sonic 1 program but it was streaming data on Sonic 2 port on converter.Shut down computer and power cycled instrument.Sonic program restarted correctly.RTS at 1608 GMT.

4/28/08

PAD

Farmer needs to plant field where instrument is located.At 1445 GMT, three technicians and I removed all soil sensors, rain gauge, and anchors.Carried 4m boom out of field.RTS at 1525 GMT.

4/21/08

PAD

IRGA SN 75B/H-0435 was returned to the CF from LiCor after repairs.Mentor request unit placed on 25m level of 60m tower and RTS.At1820 GMT, Mark Klassen and I lowered west booms, side B, of 60m tower and installed IRGA.Raised booms and started Sonic 2 and 3 programs.Program restarted correctly.Notified instrument mentor.RTS at 1850 GMT.

2/28/08

PAD

4M IRGA is showing 0 water vapor.Mentor request 4M IRGA removed and sent to LiCor for repairs.Mentor also request 25M IRGA removed from 60M tower and placed into service on 4M tower.At 1925 GMT, Ken Teske and I R/R *4M IRGA.We installed IRGA SN, on 4M tower and restarted Sonic 1 program.RTS at 1955 GMT.††

*Old- SN 75B/H-0435†† *New - 75B/H-0157

2/28/08

PAD

Mentor request *25M IRGA removed and placed on 4M system that was removed for repairs.Ken Teske and I stopped Sonic 2 program at 1853 and lowered tower.Removed 25M IRGA.RTS at 1923 GMT.

*Old - 75 B/H-0157†† *New -

2/26/08

PAD

4M IRGA, SN 0435, is not measuring water vapor.Mentor request configuration file downloaded and sent to LiCor to check.Stopped Sonic 1 program and connected laptop to IRGA.As soon as laptop was connected and error was displayed:Configuration file sent from LI-7500 contains errors.Downloaded file and sent to LiCor and mentor.Troubleshooting at this time.RTS at 2014 GMT.

10/29/07

PAD

Rodents chewed comm wires between PGS shed and 4m CO2 flux system.Repairs were made to wires on 10/29/07 but comms were not established due to time restraints.At 1842 GMT, Mark Klassen and I T/S comm issue and replaced LAN to 9 pin adapter located in comm box at base of 4m system.Comms are not working correctly.Notified mentor.RTS at 1857 GMT.

10/26/07

PAD

On 10/25/07, could not connect with 4m CO2 flux data logger.Found rodents had gotten into junction box located at base of 4m system and damaged comm cables.At 1749 GMT, Mark Klassen and I fixed several damaged comm cables for 4m datalogger, IRGA, and Sonic inside junction box at base of 4m system and between box and post where IRGA and Sonic are located.Replaced conduit between box and base of pole where instruments are located.IRGA and Sonic restarted correctly.T/S datalogger comms at this time.RTS at 2029 GMT.

10/3/07

PAD

Instrument had been removed and sent for repairs and calibration.Instrument needs to be RTS.Installed 4m *IRGA at 1736 GMT.Installed P2 Sonic and IRGA by 4M system for comparison.RTS at 2036 GMT.*New - 75B/H-0435

10/3/07

PAD

Instrument had been removed for calibration on 10/1/07 and needs to be RTS.Lowered West side B of tower at 1740 GMT.Installed 60M IRGA, SN-75B/H-0128 and raised tower.RTS at 1830 GMT.

10/3/07

PAD

Instrument had been removed for calibration on 10/1/07 and needs to be RTS.Lowered West side B of tower at 1740 GMT.Installed 25M IRGA, SN-75B/H-0157 and raised tower.RTS at 1830 GMT.

10/1/07

PAD

Mentor request 60M IRGA removed and calibrated.Lowered West side B tower at 1726 GMT and removed 60M IRGA, SN 75 B/H-0128.Mark Klassen started calibration on IRGA. Tower was raised and RTS at 1806 GMT.

10/1/07

PAD

Mentor request 25M IRGA removed and calibrated.Lowered West side B tower at 1726 GMT and removed 25M IRGA, SN 75 B/H-0157.Mark Klassen started calibration on IRGA. Tower was raised and RTS at 1806 GMT.

9/25/07

PAD

Notified by mentor 4m CO2 IRGA not registering any water vapor.Mentor request troubleshooting and removal for repairs if necessary.Connected laptop directly to LiCor LI- 7500 to monitor.Sensor shows -179.7 H20 mmol/m3.Power cycled instrument and it came back with same values.Instrument also registers 370.89 kPa.Removed *IRGA sensor and control box to send to LiCor for repairs.Notified instrument mentor.RTS at 1844 GMT.*Old - 75 B/H-0435

9/7/07

PAD

Farmer has finished baling hay in field 4m radiometer boom is located.4m boom needs to be placed back in position and connected with soil sensors.At 1500 GMT, moved 4m radiometer boom back to position in field and connected all soil sensors.Verified sensors are showing readings.Notified instrument mentor.RTS at 1530 GMT.

8/21/07

PAD

Farmer is planning on haying field where 4m CO2 flux system is located.Radiometer boom needs to be moved.At 1830 GMT, Chris Martin, Mark Klassen and I disconnected soil sensors and rain gauge and moved 4m radiometer boom out of field.Flagged soil sensors with posts and caution tape.RTS at 1915 GMT.

7/26/07

CAM

West booms need to be lowered for 25m and 60m T/H calibrations.Lowered booms at 1354 GMT.Raised booms and RTS at 1900 GMT.

4/5/07

PAD

60m instrument removed on 4/2/07 for calibration.Mentor request instrument RTS.Mark Klassen and I lowered side B, west side 60M boom and installed IRGA, SN 75B/H 0128, formerly on 25m boom on 60m boom.Tower was lowered again troubleshooting power issue with 25M IRGA.†† RTS at 1930 GMT.

4/5/07

PAD

25m instrument removed on 4/2/07 for calibration.Mentor request instrument RTS.Mark Klassen and I lowered side B, west side 25m boom and installed IRGA, SN 75B/H 0157, formerly on 4m system on 25m boom.Troubleshot instrument power issue until 1930 GMT.RTS at 1930 GMT.

4/4/07

PAD

4m instrument was removed on 4/2/07 for calibration.Mentor request instrument RTS.Installed LiCor LI-7500 IRGA, SN 75B/H 0435, formerly on 60m tower on 4m tower.Restarted Sonic1 program.RTS at 2010 GMT.

4/2/07

PAD

Mentor request 60m IRGA removed for calibration.Stopped Sonic 3 program at 1428 GMT.†† Mark Klassen and I lowered West, side B booms, and removed 60M IRGA, SN 75 B/H 0128, for calibration.

4/2/07

PAD

Mentor request 25m IRGA removed for calibration.Stopped Sonic 2 program at 1428 GMT.†† Mark Klassen and I lowered West, side B booms, and removed 25M IRGA, SN 75 B/H 0128, for calibration.

4/2/07

PAD

Mentor request 4m IRGA removed for calibration.Stopped Sonic 1 program at 1428 GMT.†† Mark Klassen and I removed IRGA, SN 75 B/H 0157, for calibration.

3/6/07

MK

Moved the 4M CO2 flux system back to its original position and reconnected all soil sensors. RTS at 2030 GMT.

3/2/07

MK

4m CO2 needs to be removed from field so farmer can spray.Moved 4m CO2 at 2210 GMT.

3/2/07

PAD

Farmer notified CF he intends to fertilize field 4m CO2 flux system is located.Disconnected all soil moisture, soil temperature, and HF plates from 4m boom.Moved 4m radiometer boom to edge of field and flagged soil sensors with caution tape for farmer.Notified instrument mentor.RTS at 2140 GMT.

2/14/07

PAD

Mentor request Kipp and Zonen net radiometer that was returned to CF from calibration installed on 4m CO2 flux system.At 2030 GMT, installed *net radiometer.Notified instrument mentor of install.RTS at 2050 GMT.††† *New - 990418 (model NR Lite, sensitivity 13.9 uV/W/m2).

12/12/06

PAD

Instrument mentor requested *Kipp and Zonen NR Lite radiometer sensor removed from 4m system and sent for calibration.Removed sensor at 2125 GMT.Took sensor to shipping and receiving and notified mentor.RTS at 2145 GMT.*Old - 990418

11/28/06

PAD

Kipp and Zonen net radiometer, (model CNR1, SN 990210), has been returned from being repaired.Mentor request instrument RTS on 4m eddy flux system.At 2003 GMT, Mark Klassen and I installed radiometer on 4m boom.Notified Sebastien Biraud of installation.Sebastien changed calibration factor of the radiometer in the datalogger program using Remote Admin.†† Instrument is displaying values on datalogger screen.RTS at 2023 GMT.

11/14/06

PAD

Farmer has planted wheat in field.4m radiometer boom and soil sensors need to be RTS.At 1936 GMT, we moved 4m radiometer boom back into field and aligned East and West.Dug two trenches and took 8 soil cores, 4 each at 10 and 20cm.Installed soil moisture sensors at 10 and 20 cm depth, soil temperature sensors at 5,15,and 25 cm depth, and soil heat flux plates at 5 cm depth.RTS at 2136 GMT.

11/13/06

PAD

4m CO2 Flux radiometer boom was removed from field on 11/9/06 so farmer could harvest soybean crop.Farmer is going to plant wheat in field today and soil probes need to be removed.At 1625 GMT, Mark Klassen and I removed soil temperature, soil moisture, and heat flux plates from field.Two of the heat flux plate wires have been cut by rodents and need to be repaired.

11/9/06

PAD

Farmer is harvesting soybeans in area.4m radiometer boom needs to be removed to allow farmer to harvest beans.Moved 4m radiometer boom out of field at 2022 GMT.Left soil sensors in ground.Tomorrow is a holiday.Radiometer boom will be returned to field as soon as possible next week.

10/11/06

CAM

West booms need to be lowered for 25m and 60m T/H calibrations.Lowered booms at 1505 GMT.Raised booms and RTS at 2117 GMT.

10/6/06

PAD

Mentor requested 4m IRGA removed and calibrated.Mark Klassen and I stopped Sonic 3 program at 1451 GMT.Removed 4m IRGA and calibrated instrument.RTS at 1611 GMT.

10/5/06

PAD

60M IRGA needs to be calibrated.Mark Klassen and I stopped Sonic 3 at 1745 GMT.Lowered side B (West side boom) of tower and removed 60M IRGA.Took IRGA to GIF trailer and calibrated.Installed IRGA, raised boom, and restarted Sonic 3 program.RTS at 2035 GMT.

10/5/06

PAD

25M IRGA needs to be calibrated.Mark Klassen and I stopped Sonic 2 at 1745 GMT.Lowered side B (West side boom) of tower and removed 25M IRGA.Took IRGA to GIF trailer and calibrated.Installed IRGA, raised boom, and restarted Sonic 2 program.RTS at 2035 GMT.

9/14/06

PAD

4m radiometer boom needs to be RTS after farmer sprayed field.At 1810 GMT, moved 4m boom back into position above soil sensors.Plugged in soil moisture, soil temperature, soil heat flux plates, and rain gauge.All sensors are showing values on display, RTS at 1830 GMT.

9/13/06

PAD

4m radiometer boom needs to be moved to allow farmer to spray field.At 1442 GMT, unplugged soil moisture, soil temperature, soil heat flux plates, and rain gauge from 4m radiometer boom and moved boom out of field.RTS at 1502 GMT.

9/11/06

PAD

Sebastien Biraud notified me that soil heat flux plate #2 was reading -200 on 4m system.At 1905 GMT, found heat flux plates #1 and #2 had been dug up by an animal.No obvious damage found on sensors.†† Reburied sensors and readings for #2 were returning to a normal value.RTS at 1920 GMT.

9/8/06

PAD

4m CO2 flux heat flux plates were removed on 9/7/06 for R/R of plate #4.New sensor was installed and heat flux plates need to be placed into service.At 1850 GMT, Mark Klassen and I buried soil heat flux plates.Changed multiplier for plate #4 from 42 to 34.6 in program.Reloaded data logger with new program.New sensor SN is H063119.RTS at 1940 GMT.

9/7/06

PAD

4m CO2 flux system shows -9s for heat flux plate #4.Mentor request heat flux plates removed #4 heat flux plate R/R.Removed heat flux plates at 1959 GMT and took to lab for repairs.RTS at 2014 GMT.

7/24/06

PAD

4m CO2 flux system heat flux plates were removed on 7/24/06 to repair broken sensor wire.Troubleshooting found bad sensor.†† Mentor requested remaining 3 heat flux plates installed until new sensor is available.At 1914 GMT, Mark Klassen and I installed heat flux plates under 4m CO2 radiometer boom at a depth of 5 cm.RTS at 1929 GMT.

7/19/06

PAD

Instrument mentor request Kipp and Zonen *net radiometer removed for calibration from 4m CO2 flux system.At 2042 GMT, Sebastien Biraud and I removed radiometer model CNR1 from 4m radiometer boom.Radiometer will be sent to Sebastien at LBNL.RTS at 2057 GMT.*Old - 990210

7/19/06

PAD

#4 heat flux plate on 4m CO2 flux systems has -9999 value.Found rodent damaged wire on #4 heat flux plate.Removed heat flux plates to repair in lab.RTS at 1940 GMT.

7/12/06

PAD

Farmer has planted field 8, where 4m CO2 Flux system is located to soybeans.4m boom needs to be moved back into field and all soil sensors need to be installed.4m boom was moved into field on 7/11/06 at 1815 GMT.At 1415 GMT, Mark Klassen and I took soil cores, installed soil moisture sensors, soil temperature sensors, and heat flux plates.Notified instrument mentor.RTS at 1615 GMT.

6/8/06

PAD

4m CO2 flux radiometer boom was removed on 6/6/06.Soil sensors need to be removed to allow farmer to harvest wheat crop.At 1930 GMT, Mark Klassen and I removed all soil moisture probes, soil temperature probes, and heat flux plates.RTS at 2010 GMT.

6/6/06

PAD

4m CO2 flux radiometer boom needs to be moved.Farmer is spraying weeds in surrounding field.At 2023 GMT, disconnected all soil sensors and rain gauge.With help, moved the 4m radiometer boom out of the field.Soil sensors will be removed tomorrow.RTS at 2038 GMT.

5/22/06

MK

Calibration completed on 25M and 60M LI-7500 units. Units are ready for reinstallation on tower. Lowered tower at 2001 GMT, installed 25M and 60M IRGA's. Raised tower at 2040 GMT. Restarted 60M software with no problems. The 25M software would not start, attempted to communicate with IRGA via laptop computer using hyper terminal. Error message (Error(received True)) was displayed several times. Power cycled 25M power twice, still no response. On 5/24/06 lowered tower, hooked up power and communications cables. Unable to communicate with IRGA, pushed reset inside IRGA enclosure, still no response. Checked all bit drivers, fiber optics, cabling and connectors. Found bad wire connection on the jumper wire installed from CTS to RTS on the serial interface connector. Fixed connection, all communications with 25M were restored, restarted software. Raised tower at 1630 GMT.

5/18/06

PAD

Mentor notified AGC value on 60m instrument has risen shortly after rain at the CF. Mentor request instrument removed from tower, inspected for possible problems or damage, calibrated and send him the calibration sheet for further instructions.Mark Klassen and I stopped Sonic 3 program at 1526 GMT.Lowered tower and removed instrument.Found large amount of bird droppings on sensor window.Cleaned optical windows, calibrated instrument and faxed calibration sheet to mentor.After mentor reviewed calibration sheets, we were instructed to change chemicals in IRGA and let them run over the weekend before calibrating again.RTS at 2026 GMT.

5/18/06

PAD

25m instrument due for calibration.Mentor request instrument removed from tower, calibrated and send him the calibration sheet for further instructions.Mark Klassen and I stopped Sonic 2 program at 1526 GMT.Lowered tower and removed instrument.Calibrated instrument and faxed calibration sheet to mentor.After mentor reviewed calibration sheets, we were instructed to change chemicals in IRGA and let them run over the weekend before calibrating again.RTS at 2026 GMT.

5/8/06

PAD

Instrument mentor request 4m IRGA changed.At 1827 GMT, Mark Klassen and I stopped Sonic 1 program.R/R *4m IRGA.Restarted Sonic 1 program.RTS at 1857 GMT.Removed unit was taken to shipping department to be sent to LBNL.*Old - 75B/H-0947*New - 75B/H-0157

4/19/06

PAD

Sebastien Biraud is on site today.Inter comparison with Ameriflux instruments is done.Request 4m radiometer boom moved to its original location.At 1450 GMT, moved 4m boom to its original location and connected soil moisture, soil temperature, soil heat flux and rain gauge to datalogger.RTS at 1530 GMT.

4/11/06

PAD

Marc Fischer is on site today.Requests 4m CO2 flux radiometer boom moved by IRGA and Sonic for inter comparison study with Ameriflux instruments. At 1549 GMT, disconnected soil temperature, soil moisture, soil heat flux plates and rain gauge.Moved 4m boom to location by IRGA.RTS at 1619 GMT.

3/9/06

PAD

CO2 values for 4m IRGA have been lower than expected.Mentor request laptop connected to 4m IRGA and values monitored.At 1900 GMT, connected laptop to 4m IRGA and called mentor.IRGA shows values of 17.58 CO2 mmol/ m3 and 427.53 umol/ mol.Values are fluctuating some.Weather conditions are light rain and windy, 7.5 m/s.Connected IRGA back to CO2 flux computer.Power cycled 4m CO2 flux system and fiber mux box located in PGS shed.Power cycled 4m media converter box located in PGS shed.Restarted Sonic 1 program.Program restarted correctly.RTS at 2000 GMT.

2/16/06

MK

Mentor has requested that the 4M IRGA should be changed out with a new one. Stopped 4M program at 1840 GMT, R/R *IRGA at 1855 GMT. Checked IRGA values using laptop computer, all values are reporting accurate. Restarted 4M program at 1908 GMT. RTS at 1910 GMT.*Old - 75B/H- 0157†† *New - 75B/H -0947

1/13/06

PAD

Mentor is troubleshooting problem with K&Z radiometer.Mentor requested wires from radiometer at AM25T multiplexer changed.At 1924 GMT, switched wires 1H-2H, and 1L-2L at AM25T multiplexer.Finished wiring change at 1930 GMT.Notified instrument mentor.RTS at 1939 GMT.

1/12/06

PAD

Removed 4m IRGA, SN 75B-0157, on 1/11/06 for calibration.At 1900 GMT, Mark Klassen and I calibrated IRGA and placed instrument back in service on 4m system.Restarted instrument at 2030 GMT.RTS at 2045 GMT.

1/11/06

PAD

Mentor request 4m IRGA removed and calibrated.Stopped Sonic 1 program at 1521 GMT.Removed LiCor IRGA SN 0157 and took to GIF trailer to calibrate.Mark Klassen and I calibrated IRGA and then changed chemicals in sensor head.Will let instrument run inside overnight and recalibrate tomorrow before returning to service.

1/10/06

PAD

CO2met computer had gone down sometime over the weekend.Mentor request computer restarted.Cooling fan is not working in laptop.Restarted computer at 1555 GMT.RTS at 1610 GMT.

1/3/06

PAD

Mentor reports CO2met computer went down sometime on 12/31/05.Restarted computer at 1642 GMT.RTS at 1657 GMT.

12/20/05

PAD

Mentor suspects problems with either K&Z CNR1 radiometer or the AM25T multiplexer.Mentor request channel 1 and 2 wires switched from radiometer at multiplexer.At 2132 GMT, switched channel 1H and 1L with channel 2H and 2L wires at multiplexer.Finished wiring at 2138 GMT.Notified instrument mentor.RTS at 2147 GMT.

12/20/05

PAD

4m system was moved over growing wheat on 12/06/05.#4 heat flux plate wire had been cut by rodents.Removed HF plate wires so repairs could be made.Mark Klassen fixed broken HF wire.At 2111 GMT, installed soil HF plates at 5cm depth under 4m system.RTS at 2131 GMT.

12/20/05

PAD

4m system was moved over growing wheat on 12/16/05.Mentor request soil temperature probes changed at datalogger.At 2111 GMT, changed soil temperature probes at datalogger as instructed.Probes #1 and #4 are now 20cm, #2 and #5 are 10cm, and #3 and #6 are 5cm. depth.

12/16/05

PAD

Farmer has planted field CO2 flux system is located in to wheat.Farmer planted around 4m CO2 flux system.Mentor requested system moved so it is located over growing wheat.At 1900 GMT, Zack Waddell and I removed soil moisture probes, soil temperature probes, rain gauge, and heat flux plates.Moved 4m system over growing crop.Took soil cores from 10cm and 20cm depths.Installed soil moisture, soil temperature probes, and rain gauge.One heat flux plate wire was cut by rodents.Will install HF plates after repairs are made.RTS at 2200 GMT.

12/12/05

PAD

Instrument mentor emailed he could not connect with CO2met computer.Found computer was down.Previous troubleshooting found computer fan not working.Restarted computer.RTS at 1913 GMT.

12/8/05

PAD

Found CO2met computer down.Restarted computer at 1648 GMT.RTS at 1703 GMT.

12/7/05

PAD

Sebastien Biraud notified he couldnít remotely access PGS, CO2met, CO2flux, or CO computers in PGS shed.†††† Reset Allied Telesyn media converters located in PGS shed and AOS trailer.Computers can now access network.Notified instrument mentor.RTS at 1735 GMT.

12/5/05

PAD

Found CO2 met computer down.Previous troubleshooting found cooling fan not working.Computer has picture from 12/1/05 at 1539 GMT.Don't know what time computer shut down.Restarted computer at 1528 GMT.RTS at 1543 GMT.

11/21/05

PAD

CO2 met computer is down.Previous troubleshooting found computer cooling fan is not working.Restarted computer at 1705 GMT.RTS at 1720 GMT.

11/7/05

PAD

CO2 met computer is down.Previous troubleshooting shows computer cooling fan is not working.Restarted COmet computer at 1543 GMT.RTS at 1558 GMT.

10/25/05

PAD

CO2 met computer is down.Previous troubleshooting found computer fan is not working.Restarted computer and checked that computer will connect with data logger and network.RTS at 1633 GMT.

10/11/05

PAD

Mentor notified CO2 met computer is down again.Started computer at 1421 GMT.Computer will connect with data logger and network.RTS at 1436 GMT.

10/4/05

PAD

At 1855 GMT observed Sonic 1 from 4m CO2 flux system will not graph on computer display.Stopped and restarted Sonic 1 program.Segmentation fault error came up.Power cycled 4m IRGA and restarted program.Sonic 1 program restarted correctly and will now display graph on computer.RTS at 1910 GMT.

10/3/05

PAD

Mentor notified he couldnít connect with 4m CO2 flux system data logger using CO2met computer.Power and communication cables were removed from 4m CO2flux system on 9/30/05 to allow farmer to harvest corn in field instruments are located in.At 1925 GMT, plugged power and communication cables back into 4m CO2flux system.CO2met computer will now connect with data logger.RTS at 1945 GMT.

9/30/05

PAD

Farmer is going to harvest corn that is planted in field 8 where 4m CO2flux system is located.System needs to be removed.At 1740 GMT, started removing soil moisture probes # 2 and # 4.Decided there is not enough corn under 4m boom to justify moving system.Reinstalled soil moisture sensors # 2 and # 4.Unplugged communication and power cables and pulled them back out of the way of combine.RTS at 1800 GMT.

9/13/05

PAD

Mentor notified 4m Sonic 1 is not collecting data.At 1747 GMT, checked Sonic 1 program.Graph would not display.Exited Sonic 1 program and restarted.Program tried to restart but shows a Segmentation fault.Connected 4m IRGA to laptop but cannot connect to instrument.Power cycled instrument and started Sonic 1 program.Sonic 1 data graph will now display.Notified instrument mentor.RTS at 1802 GMT.

8/26/05

PAD

Mentor notified he cannot remotely access 4m CO2 met computer.Computer was down.Restarted computer and unplugged power.Computer battery shows 100%.After 5 minutes, battery shows 95%.Computer is on UPS power and UPS passes all tests.RTS at 1526 GMT.

8/15/05

PAD

4m CO2 flux Sonic 1 program will not graph.Stopped and restarted program.Display shows segmentation fault.Power cycled instrument and restarted Sonic 1 program.RTS at 1945 GMT.

8/8/05

PAD

Instrument mentor informed he cannot remotely access CO2 met computer.At 1740 GMT, found computer was not running.Checked power supply for computer.Started computer and checked computer could access network.Notified instrument mentor.RTS at 1755 GMT.

8/5/05

CAM

Need to stop the Sonic programs for tower maintenance.Stopped 4, 25, and 60M Sonic programs at 1328 GMT.Restarted programs following maintenance at 1700 GMT.

8/4/05

CAM

Need to stop the 25 and 60M Sonic programs for tower maintenance.Stopped 25 and 60M Sonic programs at 1327 GMT.Restarted programs following maintenance at 2030 GMT.

7/29/05

PAD

Instrument mentor request 60M IRGA placed in service.Shut down 25M Sonic 2 program at 1843 GMT.Lowered West booms and installed *60M IRGA.Set instrument parameters as per mentor's instructions.Raised tower and RTS at 2013 GMT.*New - 0435

6/23/05

PAD

Instrument mentor requests 4m CO2flux rain gauge raised to a height above corn plants in field # 8.At 1925 GMT, installed conduit in ground and attached rain gauge to taller conduit.RTS at 1940 GMT.

6/23/05

PAD

Heat flux sensors were removed on 6/16/05 to repair damage caused by rodents. Mike Rainwater repaired heat flux wires as needed.Ordered conduit to be placed over heat flux wires to protect them.Placed wires in conduit from sensor pad up to where extra wire is coiled at data logger box.On 6/23/05, installed heat flux plates at 1355 GMT.RTS at 1415 GMT.

6/16/05

PAD

Sebastien Biraud notified heat flux plate #4 is not showing a value.Found wire had been cut by rodents.This has been an ongoing problem, as wires are not protected by conduit.Removed all heat flux wires at 1548 GMT.Mike Rainwater splices wires as needed.Had enough flexible plastic conduit for one wire.Ordered extra conduit.Will install heat flux wires when repairs are complete.RTS at 1600 GMT.

6/13/05

PAD

Unable to communicate with 60m IRGA.Need to lower tower to troubleshoot communication problem.Lowered side B, 25m and 60m booms at 1748 GMT.Found IRGA control box was not latched.Control box has water in it from recent rains.†† Removed instrument and control box.Contacted instrument mentor.RTS at 1838 GMT.

6/10/05

PAD

On 6/9/05, 60m IRGA instrument stopped transmitting data.Instrument mentor requests CO2flux computer rebooted.Rebooted CO2flux computer at 1921 GMT.60m IRGA will still not communicate with computer.4m and 25m Sonic and IRGA instruments restarted correctly.Notified instrument mentor.Troubleshooting at this time.RTS at 1936 GMT.

6/9/05

PAD

At 1912 GMT, 60m Sonic 3 instrument would not graph on computer.Restarted program but instrument would not graph.Mike Rainwater and I power cycled instrument, checked bit drivers and tried to communicate with instrument with laptop.Laptop would not connect with LiCor software but we could connect with hyperterminal program.Instrument does not appear to be sending any usable data.Contacted instrument mentor.Waiting on reply.RTS at 2010 GMT.

6/9/05

PAD

At 1930 GMT, 25m CO2 flux system would not graph on computer.Power cycled instrument and restarted program.Reseated tower communications and power connectors at 2006 GMT.Restarted 25m CO2 flux system.RTS at 2010 GMT.

5/27/05

PAD

Sebastien Biraud notified that 4m Sonic was not operating correctly.Sonic 1 would graph but some CO2L and H20L values were not changing.Stopped and restarted program.Power cycled instrument at tower and restarted program.Program restarted correctly and values are now changing.RTS at 2033 GMT.

5/20/05

PAD

Farmer needs to spray field 8 for weeds where 4m CO2 flux system is located.At 1435 GMT, Mike Rainwater, James Martin, Mark Klassen, and I unplugged soil moisture, soil temperature, and heat flux sensors from 4m tower.Moved tower out of the way.Notified instrument mentor and farmer.RTS at 1455 GMT.

4/18/05

PAD

Soil sensors need to be installed on 4m CO2 flux system.Modified soil moisture sensors as instructed by mentor.Mike Rainwater and I installed soil moisture, soil temperature and heat flux plates under 4m tower.RTS at 2030 GMT.

4/15/05

PAD

Instrument mentor requested 4m CO2 flux system relocated to new boom 25 m south of current boom.Upgrade started on 4/13/05.At 1330 GMT, Mike Rainwater, Dan Nelson and I installed leg extensions on 4m boom, installed radiometer, and wind sensors to new tower.†† All instruments installed are showing values on computer display.RTS at 1900 GMT.

4/14/05

PAD

Instrument mentor requested 4m CO2 flux system relocated to new boom 25 m south of current boom.At 1500 GMT, Mike Rainwater and I connected extensions to power and communication cables, set up new 4m boom and moved data logger box and radiometers.

4/12/05

PAD

Sebastien Biraud notified CO2 flux computer is not collecting data from instruments.Instruments appear to be working and trying to communicate with computer.Stopped CO2 flux data graphs.Computer would not bring up new terminal to restart program.Rebooted computer.†† All CO2 flux instruments will now graph on computer.Notified instrument mentor.

3/30/05

PAD

Need to remove soil sensors from 4m CO2flux system so farmer can plant corn.At 2121 GMT, Mike Rainwater and I removed soil moisture sensors, soil temperature sensors, and heat flux plates.RTS at 2151 GMT.

3/28/05

TMG

Need to install RAdmin on the CO2met computer. Stopped PC208W software @ 1513 GMT and installed RAdmin @ 1535 GMT. Rebooted the computer @ 1543 GMT and RTS @ 1545 GMT.

2/22/05

PAD

At 1718 GMT, found CO2 flux and CO2 met computers were off.PGS sample pumps were not running.Computer monitor did not have power.††† Found tripped breaker in PGS shed caused by shorted wire in PGS pump rack.Reset breaker and UPS for CO2 met and CO2 flux computers.Restarted computers.Power cycled 25m and 60m instruments to establish communication with computer at 1752 GMT.Restarted Sonic 2 and 3 programs.RTS at 1808 GMT.

1/7/05

PAD

Sebastien Biraud notified wind speed indicator on 4m CO2flux system was not working.Wind speed indicator was frozen due to recent ice storm.††† Removed ice from sensor.Sensor started working correctly.Notified instrument mentor.RTS at 2030 GMT.

12/29/04

PAD

Instrument mentor requested support arms added to stabilize 4m boom.At 1915 GMT, met with facilities and added support arms.Leveled instruments located on boom.RTS at 1945 GMT.

12/21/04

PAD

Instrument mentor requests Li-Cor Quantum sensor that is facing up for comparison, changed to face the ground on 4m tower.Removed Quantum sensor (SN Q34242) and repositioned sensor to face down.Took pictures of sensor location on boom as per mentor's request.RTS at 1718 GMT.

12/15/04

PAD

Soil sensors have been removed to allow farmer to plant wheat in field where 4m tower is located.Farmer has decided not to plant wheat due to delay caused by wet conditions.Instrument mentor requested soil sensors installed.At 2030 GMT, Mike Rainwater and I installed soil moisture sensors, soil temperature sensors, soil heat flux plates, and took soil cores.Made sure all sensors were active and displaying on computer.Notified instrument mentor.RTS at 2200 GMT.

10/20/04

PAD

Sebastien Biraud notified PC208 program was not running on CO2Met computer and he cannot connect with data logger at 4M tower.Started PC208 program at 1756 GMT.Computer will connect with data logger.Notified instrument mentor.RTS at 1811 GMT.

9/27/04

CAM

West booms need to be lowered for 25m and 60m T/H calibrations.Lowered booms at 1432 GMT.Raised booms and RTS at 2105 GMT.

9/9/04

PAD

During PM rounds, observed there were two Sonic 3 programs running on CO2flux computer for 60M Sonic.††† Notified instrument mentor.Stopped and restarted Sonic 3 program.Made sure program started correctly.RTS at 1625 GMT.

8/5/04

PAD

Soil sensors were removed from under 4m CO2 flux system in field 8 on 7/7/04 so farmer could till under radiometer boom.Farmer tilled field again on 8/3/04.Soil sensors need to be installed.At 1930 GMT buried soil moisture sensors at 5 and 25 cm depth.All sensors are showing values on computer.RTS at 2010 GMT.

8/4/04

PAD

Soil sensors were removed from under 4m CO2 flux system in field 8 on 7/7/04 so farmer could till under radiometer boom.Farmer tilled field again on 8/3/04.Soil sensors need to be installed.At 1929 GMT, Mike Rainwater and I took (4) 25 cm soil cores, buried soil temp sensors at 5,15, and 25 cm, and heat flux plates at 5 cm.All sensors are showing values on computer.RTS at 2049 GMT.

7/7/04

PAD

Soil sensors need to be removed from under 4m CO2flux system to allow farmer to till ground below instruments.At 1928 GMT, Mike Rainwater and I removed heat flux plates, soil moisture sensors and soil temperature sensors below radiometer boom.RTS at 2018 GMT.

7/6/04

PAD

CO2met computer will not connect with 4m data logger.Data logger lid was left open from troubleshooting previous communication problem.Left box lid open to dry out.Power cycled data logger at 1650 GMT.CO2met computer will now connect with data logger.RTS at 1705 GMT.

6/3/04

PAD

During troubleshooting 4m CO2flux system network problems, CO2flux computer was rebooted.Power cycled 25M and 60M west booms at 2030 GMT.There is no failure with this instrument.

6/3/04

PAD

During PM rounds, 4m CO2 flux system would not graph on computer display.Severe weather at site yesterday.Tried to restart Sonic program for 4m system.Power cycled hub in PGS shed.Power cycled IRGA instrument at 4m tower.Mike Rainwater and I hooked laptop to 4m IRGA and SONIC and both instruments would communicate with laptop.Power cycled port hubs located at base of 4m tower and inside PGS shed.Rebooted CO2 flux computer.All CO2 flux systems came back up and would graph on computer.RTS at 2105 GMT.

4/29/04

MR

The calibration and work has been completed on this system and it is now ready to be raised.Raised the 60M tower.RTS at 1341 GMT.

4/26/04

JEM

The IRGAís (LI-7500) need to be removed so they will be ready for Mark Fisher to calibrate on 4/27/04. Closed the program @1527 GMT.Removed the 4M IRGA @1610 GMT.†† Lowered the 25M boom to the ground.Removed the 25M IRGA @1535 GMT.Lowered the 60m boom to the ground.Removed the 60M IRGA @1547 GMT.Placed components in the GIF trailer for Mark Fisher.Booms left at ground level pending calibration of the IRGAís.

1/29/04

PAD

Instrument mentor notified problem with 60M CO2 flux system communicating with computer.Power cycled 60M instruments on West side of boom at 2209 GMT.Power cycled USB port for Sonic 2 and 3.Contacted instrument mentor and troubleshot potential causes.RTS at 2302.

1/29/04

PAD

While troubleshooting 60M CO2 Flux system, 25M CO2 flux system was shut down.Power cycled 25M CO2 Flux system at base of tower at 2209 GMT.Restarted Sonic 2 program on computer.RTS at 2224 GMT.

11/3/03

PAD

Instrument mentor notified Heat Flux 4 readings were bad.On 10/30/03, found rodents had chewed through the Heat Flux wire.On 11/3/03 at 1857 GMT, found rodents had chewed through Heat Flux 4 wire in two places.Spliced the broken wires.Will contact mentor and suggest running wires through conduit to help with rodent problem.RTS at 1932 GMT.

8/20/03

PAD

4M CO2 Flux system was not graphing data on computer.Contacted instrument mentor.Cycled power to 4M CO2 Flux system at 1720 GMT.Restarted Sonic 1 program on computer and checked that data was graphing.RTS at 1735 GMT.

8/14/03

PAD

Dave Cook on site for 25 and 60M T/H calibration.West booms lowered at 1515 GMT on 8/14.Raised booms on 8/15 and RTS at 1939 GMT.Some communication problems with 25M CO2 Flux instrument after raising booms.Cycled power to instrument at base of tower and rebooted computer.Troubleshooting with instrument mentor at this time.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.Power was restored to the Central Facility at approximately 1945 GMT.RTS @2000 GMT.

3/21/02

CAM

Check of CO2 Flux system following a hard power bump found that the computer was locked up.The power LED was flashing and the hard drive activity light was lit solid.The computer could not be power cycled with the on/off switch.Disconnected the AC power cord from the computer for several seconds until the power LEDís went off.Reconnected AC power cord and system then booted up normally.RTS at 2001 GMT.

3/4/02

CAM

Instrument mentor notified Site Ops that he was unable to connect to the CO2 Flux computer.Checked on the system and found that the UPS was not supplying power to the computer.Site power was out for 1.5 hours on 3/2, which had drained the UPS batteries and it shut down.Powered the UPS on at 1755 GMT.CO2 Flux RTS status O.K. at 1757 GMT.The UPS was displaying an alarm for low battery voltage.Will see if the alarm clears after the batteries charge longer.

2/12/02

CAM

Instrument mentor requested that the port1flux laptop be powered on.The laptop shut down during the recent power outage and does not auto-restart.Powered the laptop on at 2035 GMT.RTS at 2038 GMT.

12/3/01

CAM

Instrument mentor notified Site Ops that the CO2 system was not generating any data since 12/1.Status check found an error message on the screen indicating Com Port 1 could not be opened.Rebooted the computer at 1721 GMT.RTS status O.K. at 1723 GMT.

11/26/01

CAM

Instrument mentor notified Site Ops that the CO2 system was not generating any data.Status check found an error message on the screen indicating Com Port 1 could not be opened.Rebooted the computer at 1851 GMT.RTS status O.K. at 1853 GMT.

11/05/2001

TMG

Mentor reports that the 60m Flux has not been working properly since D4 program was installed and requests that the following steps be performed on the 60m Flux computer. 1.) Move raw and processed data to e:\Data and 2.) Disable D4 time protocol program. Technician moved all the raw data files in the collection directory (raw (M*)to the larger NTFS partition (e:\data\). Technician then disabled the D4 clock program. Computer rebooted and RTS at 2100 GMT.

10/25/01

CM/TG

The CO2 flux time was 9 minutes and 26 seconds ahead of SDS time.Installed D4 program on the CO2 computer and configured it to sync the clock to SDS every 15 minutes.Updated the time at 1835 GMT.RTS status O.K. at 1840 GMT.

10/24/01

CAM

Instrument mentor informed Site Ops that no data was being collected from the CO2 Flux system for the past couple of days and requested that the computer be rebooted.†† The program was running and the system appeared normal.Stopped the program and rebooted the computer as requested at 1743 GMT.†† When the program restarted there was an error opening Com Port 1.Rebooted again at 1745 GMT and this time no errors were encountered.RTS at 1747 GMT.††† Mentor later found that the system had recovered prior to the reboot.

7/3/01

CAM

Instrument mentor notified Site Ops that the CO2 Flux system had not collected any files since Friday, 6/29.T/S found that the Gillsonic was minimized which causes it to freeze.Maximized Gillsonic program and RTS status O.K. at 1312 GMT.Suspect that the computer was inadvertently power cycled during maintenance to the PGS system on 6/29 and the program never started.

5/23/01

MR

The instrument mentor wants the 60 meter instruments power cycled. Power cycled the instruments at 1455. RTS at 1510.

4/13/01

Mentor

The CO2 flux computer has been relocated to the new building at the base of the 60M tower. The new building was installed to house the Precision Gas Sampling (PGS) equipment.

3/21/01

CAM

Instrument mentor asked if the CO2 Flux computer could be configured to automatically restart after a power failure to minimize data loss. Stopped the CO2 Flux program at 1800 GMT. Entered the BIOS setup utility and enabled the AC Power Loss Restart option. Restarted the computer and program and then disconnected the AC power for 5 seconds to test the restart. Applied AC power and the computer powered itself on as desired. However, the Logon prompt came up and the CO2 Flux program would not start until the proper username and password were entered to log on. RTS the system at 1810 GMT to research auto logon capabilities. Found that the registry could be modified to perform an auto logon. Stopped the program at 2015 GMT and made the necessary registry changes. Tested the system several times and found that the computer will now automatically restart and log on. RTS at 2058 GMT.

2/26/01

CAM

Marc Fischer notified Site Ops that he had not received files from the instrument since 2/24. T/S found that the C02 Flux computer was powered off. The site experienced a 1 hour power outage on 2/24 which had drained the UPS batteries and caused the computer to power off. Restored computer power at 2029 GMT. RTS status O.K.

2/7/01

CM

Marc Fischer notified Site Ops that he was not receiving data from the instrument. T/S found that the PC was powered off. The site experienced some problems with one leg of AC power on 2/5. This caused some problems with some of the smaller UPS's and the UPS probably shut the computer down then. Powered the computer on and RTS at 1922 GMT.

11/17/00

KT

At 1937 GMT powered off computer to install 360 UPS. At 1941 GMT plugged instrument computer into UPS and left monitor plugged into surge protector. At 1942 GMT powered computer on. Status ok. RTS at 1952 GMT.

11/16/00

TG

Instrument mentor reported that the flux system stopped collecting data at 1930 GMT on 11/14. T/S found the computer was not logged on to. Logged on and the flux program started working. Apparently, the system had a power bump the day before that caused the computer to reboot. An operator must log into the system for the software to work properly.

11/13/00

TG

Set the CO2 Flux clock to CST time. The computer was previously set to PST.

11/9/00

CW

Rebooted CO2 Flux computer following tower maintenance.

9/20/00

TG

Mentor reported that the computer was locked up. T/S indicates that root drive was filled with data files. Technician created a data directory and copied all data files into the new directory. Rebooted computer and RTS status ok.

9/13/00

CW

CO2 Flux station installed on the 60m tower from 1730-2230 GMT.